Using the information below to create an Excel worksheet, answer the following questions. Show 2 decimal places. (This is a pWhat is the EVWPI? Answer: What is the best EMV? Answer: What is the EVPI? Answer:
Answer #1

1) Expected value of EVwPI = (Best payoff of strong)x (Probability of strong)+ (Best payoff of medium)x (Probability of medium)+ (Best payoff of weak)x (Probability of weak)

=(165 x 0.2) + (153 x 0.35) + (91 x 0.45)

=33 + 53.55 + 40.95

=127.50

2) EMV = Weighted average of the payoffs for a decision alternative

Option A = (165 x 0.2) + (88 x 0.35) + (64 x 0.45) = 92.60

Option B = (142 x 0.2) + (153 x 0.35) + (27 x 0.45) = 94.10

Option C = (76 x 0.2) + (82 x 0.35) + (91 x 0.45) = 84.85

Best EMV = 94.10

3) EVPI = EVwPI - EVwoPI (Best EMV)

= 127.50 - 94.10

= 33.40